OpenStudy (anonymous):

How would this work please help... Create your own third degree polynomial that when divided by x + 2 has a remainder of –4.

OpenStudy (anonymous):

so basically you have to create a polynomial that has (x+2) you can go with anything that is in the form of : (ax+c)(x+2)-4

OpenStudy (anonymous):

And the question has a little issue and that is the fact that in most math refrences the remiander must be positive.

OpenStudy (anonymous):

can you show me an example cause i still dont get it :/

OpenStudy (anonymous):

for example A=(x-4)(x+2)-4. think about it this way: (x-4)=b, (x+2)=q, r=-4 and here is the general rule of division: A=bq+r don`t they look similar?

OpenStudy (anonymous):

yes so i would just have to plug in numbers into this formula and my plynomial equation would be good?

OpenStudy (anonymous):

yeah in the first formula that i posted , a and c can literally be any number.( except, "a" can`t be zero because then it wont be third degree.)

OpenStudy (anonymous):

ok thank you for your help :)